OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6.

OpenStudy (anonymous):

3(x-2)=6 Does viewing it like that help?

OpenStudy (anonymous):

No, I have no clue on how to answer this question...

OpenStudy (bambimonster):

3(x-2)=6 3x-6=6 nw solve for x